OpenStudy (anonymous):

can someone just tell me what the cube of 128 is?

OpenStudy (anonymous):

Cube as in cube root?

OpenStudy (anonymous):

yes

OpenStudy (taylormarie213):

11.31

OpenStudy (nurali):

\[\sqrt[3]{128}\] \[\sqrt[3]{64*2}\] \[\sqrt[3]{2*2*32}\]

OpenStudy (wolf1728):

No I think it refers to 128*128*128 or 128³

OpenStudy (nurali):

\[\sqrt[3]{2*2*2*16}\] \[4\sqrt[3]{2}\]

OpenStudy (wolf1728):

Well did you want the cube or the cube root?

OpenStudy (anonymous):

Is it 128 cubed, or the cube root of 128

OpenStudy (anonymous):

Cube root of 128 = 5.0296842 128 cubed = 2097152

OpenStudy (wolf1728):

I calculate cube root = 5.0396841996 and cube = 2,097,152